OpenStudy (anonymous):

Find the greatest common factor of 270 and 360. (Give the answer in the numerical form in the top box and in exponential form by filling in the boxes for exponents.)

OpenStudy (anonymous):

Can someone help me on this I am stuck

TheSmartOne (thesmartone):

Hi! :D

OpenStudy (anonymous):

Hi

TheSmartOne (thesmartone):

What two numbers multiply to equal 270?

OpenStudy (anonymous):

135*2

TheSmartOne (thesmartone):

Also, welcome to OpenStudy! :D

OpenStudy (anonymous):

Thank you

TheSmartOne (thesmartone):

ok so 270 = 135*2 let's just keep factoring all the numbers until we're left with prime numbers let's divide 135 by 5. What is 135/5 = ?

OpenStudy (anonymous):

27

TheSmartOne (thesmartone):

ok so 270 = 27*5*2 Let's break down 27. What is 27/3 = ?

OpenStudy (anonymous):

its 9

TheSmartOne (thesmartone):

ok 270 = 9*3*2*5 what is 9/3 = ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

9\3 is 3

TheSmartOne (thesmartone):

ok so 270 = 3*3*3*2*5 \(\sf\Large 270 =2* 3^3*5\) Now let's factor down 360

TheSmartOne (thesmartone):

What 2 numbers multiply to 360? :)

OpenStudy (anonymous):

180*2

TheSmartOne (thesmartone):

ok, and 180 /2 = ?

OpenStudy (anonymous):

=90

TheSmartOne (thesmartone):

and 90/2 = ?

OpenStudy (anonymous):

45

TheSmartOne (thesmartone):

so far we have: 360 = 45 *2*2*2 ok 45/5 = ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

=9 sorry about that my computer just froze

TheSmartOne (thesmartone):

and 9/3 = ?

OpenStudy (anonymous):

3

TheSmartOne (thesmartone):

it's ok :)

TheSmartOne (thesmartone):

so now we have: 360= 3*3*2*2*2*5 \(\sf\Large 360 = 2^3*3^2 *5\)

TheSmartOne (thesmartone):

\(\sf\Large 270 =2* 3^3*5\) \(\sf\Large 360 = 2^3*3^2 *5\) So these are our 2 numbers factored down. :) So how many "2" are common in 270 and 360? For example, there is one 2 in 270. And there is three 2's in 360. So they only have in common one 2. How many 3's are common in between 270 and 360? :)

OpenStudy (anonymous):

there is 4 3s

TheSmartOne (thesmartone):

mhmm, not sure how you got 4 we have 3*3*3 in 270 and 3*3 in 360 therefore, they only have two 3's in common

OpenStudy (anonymous):

oh my bad

TheSmartOne (thesmartone):

and we have one 5 in common, right? :)

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

so what is 2 x 3 x 3 x 5 = ?

TheSmartOne (thesmartone):

that is the same thing as 2*3*3*5 = ?

OpenStudy (anonymous):

its 90

TheSmartOne (thesmartone):

and that's your GCF of 270 and 360 :)

TheSmartOne (thesmartone):

there was another simple way you factor 10 out of both 360 and 270 you're left with 36 and 27 then you see that the GCF of 36 and 27 is 9 and then you multiply 10 to everything so the GCF of 360 and 270 is 90 :P

OpenStudy (anonymous):

:) well thx so much for your help

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @tryingtobeateacher @~

TheSmartOne (thesmartone):

Anytime! :D